OpenStudy (anonymous):

xy+xz/xz-z^2 xy-yz/xz+yz%y^2+yz/x^2+xy

OpenStudy (anonymous):

Perform the indicated operation

OpenStudy (anonymous):

@coral

OpenStudy (highlyeducated87):

Answer: -xyz^2+2xy+yz/x^2-y/x+z/y+1

OpenStudy (anonymous):

i think xy+xz/xz-z^2 * xy-yz/xz+yz / y^2+yz/x^2+xy = x(y+z)/z(x-z) * y(x-z)/z(x+y) x x(x+y)/y(y+z) = x(y+z).y(x-z).x(x+y) / z(x-z).z(x+y).y(y+z) = x^2 / z^2

OpenStudy (highlyeducated87):

It may be but i did the work and got that answer